Circular No. 6833

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

S/1997 U 1 This ephemeris, from the orbit given on IAUC 6780 for the fainter of

the two new Uranian satellites, is likely to be in error by several arcmin. The columns headed Δα′′ and Δδ′′ give the offsets from Uranus in α (including the cos δ factor) and δ, each in arcsec, the sign being at the end.

1998 TT α_{2000} δ_{2000} Δ r_{U} ϵ Δα′′ Δδ′′
Mar. 8 20^{h}53.^{m}52 −18^{o}12.^{′}7 20.666 0.061 36.^{o}5 358− 418−
18 20 55.40 −18 05.2 20.554 0.061 46.0 388− 422−
28 20 57.07 −17 58.5 20.421 0.061 55.5 415− 423−
Apr. 7 20 58.48 −17 52.9 20.272 0.062 65.1 440− 422−
17 20 59.60 −17 48.4 20.112 0.062 74.6 463− 419−
27 21 00.42 −17 45.1 19.944 0.062 84.1 484− 414−
May 7 21 00.92 −17 43.2 19.773 0.062 93.7 502− 406−
17 21 01.09 −17 42.6 19.603 0.062 103.3 518− 395−
27 21 00.93 −17 43.4 19.441 0.062 113.0 532− 381−
June 6 21 00.46 −17 45.4 19.289 0.062 122.7 542− 365−
16 20 59.69 −17 48.7 19.154 0.062 132.4 550− 346−
26 20 58.65 −17 52.9 19.037 0.062 142.2 556− 325−
July 6 20 57.39 −17 58.0 18.944 0.062 152.1 557− 301−
16 20 55.96 −18 03.7 18.877 0.061 162.0 556− 275−
26 20 54.41 −18 09.7 18.838 0.061 171.9 551− 247−

S/1997 U 2 P. D. Nicholson, Cornell University, reports that B. J. Gladman and

J. A. Burns have located images of the brighter Uranian satellite (IAUC 6764, 6780, etc.) on plates taken with the Canada-France-Hawaii Telescope in 1984. Semiaccurate measurements (uncertainty ±3′′) for them are listed below, together with some additional recent CCD data:

UT  α_{2000}  δ_{2000}  Observer
1984 June 1.5556  16^{h}40^{m}44.^{s}2  −22^{o}07′02′′  Cruikshank
2.5410  16 40 34.0  −22 06 40  _{”}
1997 Nov. 30.06641 20 32 40.90 −19 22 30.3 Offutt
30.09408 20 32 41.08 −19 22 29.6 Holman
Dec. 17.06851 20 35 43.90 −19 11 15.5 Offutt

D. P. Cruikshank (Mauna Kea). Measurer B. J. Gladman. W. B. Offutt (Cloudcroft). 0.6-m f/7 Ritchey-Chrétien. M. J. Holman and W. R. Brown (Mt. Hopkins). 1.2-m reflector.

1998 March 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden